Direction: Answer the following problems stated below. Show your solutions.



1. Determine the amount of current produced by a charge of 4.55x10^-17 C 5s along a straight conductor.



2. How many electrons will flow through an appliance for each if the current through the appliance is 0.85A



3. Consider a current of 0.55A flowing steadily through a conductor. How many charges pass through the wire in a span of 2 minutes?

1
Expert's answer
2022-02-07T08:42:59-0500

1.

"I=\frac{q}{t}=\rm\frac{4.55*10^{-17}\: C}{5\: s}=9.1*10^{-18}\: A"

2.

"q=It=\rm0.85\: A*5\: s=4.25\: C"

"N=\frac{q}{e}=\rm\frac{4.25\: C}{1.6*10^{-19}\: C}=2.66*10^{19}"

3.

"q=It=\rm0.55\: A*2*60\: s=66\: C"